OpenStudy (anonymous):

Please help and show all work: Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

Y=4x+7 y-(-3)= 4x-8 y+3=4x-8 y=4x-11

OpenStudy (anonymous):

y-y1=m (x-x1) equation look familiar?

OpenStudy (anonymous):

uhm that is not the answer... m=-(1/4) and not 4... parallel lines have the same gradient, but perpendicular lines have gradients whose product is -1. so the answer is y=(-1/4)x -(5/2)